OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 3. Please explain in sentences how you arrived at the answer. Thanks

OpenStudy (anonymous):

first you need to find \[\log_{16}(32)\]

OpenStudy (anonymous):

you have a choice, math teacher way or cheat way \[16^y=32\] \[2^{4y}=2^{5}\] \[4y=5\] \[y=\frac{5}{4}\]

OpenStudy (anonymous):

then solve \[\frac{5}{4}=x+3\] in one step (subtract 3) and you are done